# Spin Theory Based on the Extended Least Action Principle and Information Metrics: Quantization, Entanglement, and Bell Test With Time Delay

### v2

@@ III.1 Spin Model
In order to explain the existence of electron spin, additional assumptions on electron properties are needed. Historically, several models have been proposed to explain the existence of spin. The most popular model is to consider the electron as a rotating rigid body with uniform distributed charge. Denote the electron magnetic moment as $\vec{\mu}$ and the angular momentum as $\vec{L}$. The relation between $\vec{\mu}$ and $\vec{L}$ can be derived if we consider the electron as a localized electrical distribution with current density $\vec{j}(\vec{r})$, and the magnetic moment of electron is the magnetic dipole moment, which is calculated as
@@
where $\vec{p}(\vec{r})$ is the momentum density. However, one of the problems for this model is that when the radius of the rigid body is sufficiently small, the edge is moving faster than the speed of light Tomonaga . There is also a simpler model that considers an electron as a point particle with intrinsic magnetic moment and intrinsic angular momentum and they still satisfy the relation (3). Such a model offers no explanation on how the intrinsic angular moment or intrinsic magnetic moment originate.
Built upon the success of applying the extended least action principle in deriving quantum theories Yang2023 ; Yang2024 , we wish to explain the origin of the intrinsic magnetic moment and angular momentum by applying the same principle but adding additional refinement on the vacuum fluctuations assumption in Section II. The electron exhibits random displacement due to vacuum fluctuation and such displacement is described as a vector. The displacement vector is intrinsically local in the sense that it is independent of the reference of origin of the coordinate system. If we further assume the displacement vector contains not only translational components, but also rotational components, then the electron will exhibit intrinsic angular momentum due to such random rotation. In Appendix B, using the extended least action principle, we show that if the rotation of the displacement vector is a circular movement, the averaged magnitude of intrinsic angular momentum is $L_{s}=\hbar/2$, while the orientations of the intrinsic intrinsic angular momentum is completely random. That is, the intrinsic angular momentum is isotropic and its orientation is randomly fluctuating. Although our model is very different from the rigid body model, we assume that the relation (3) still holds except with a factor $g_{s}$ to be determined further
@@
Since the orientation of the magnetic moment is always opposite to the orientation of intrinsic angular momentum, the orientation of the intrinsic magnetic moment is randomly fluctuating as well. Although the result in Appendix B is impressive enough, the assumption of circular motion of displacement vectors due to vacuum fluctuation is rather strong. Instead, we will just assume the existence of intrinsic angular momentum with magnitude of $\hbar/2$ and with random orientation for an electron. To summarize, in addition to Assumption 1 and 2 as described in Section II, we need
@@
As to be shown next, with this additional assumptions and applying the extended least action principle, we can recover the predictions of quantum theory of the electron spin, including the discreteness of spin measurement results, the Pauli-Schrodinger equation, and entanglement between two correlated electron spins [^1] It is desirable to develop a better model than that described in Appendix B to derive the $\hbar/2$ magnitude of intrinsic angular momentum, but we leave it as a future research topic..
Suppose we choose a reference frame such that the electron is at rest, that is, the average translational momentum is zero. The electron still exhibits an intrinsic angular momentum with the random orientation according to Assumption 3. We want to derive the probability distribution of the intrinsic angular momentum orientations. If there is no external magnetic field applied, the probability distribution is simply an uniform distribution according to Assumption 3. Now, suppose an external magnetic field along the direction of the $z$-axis, $B_{z}$, is applied, the probability distribution of the magnetic moment orientations is no longer uniform. We show next that this probability distribution can be derived from the extended least action principle.
Due to the interaction of electron magnetic moment and the external magnetic field, the electron is experiencing the Larmor precession around the $z$-axis, as shown in Figure 1. Denote the probability density of the intrinsic angular momentum orientations as $p(\theta,\varphi)$ where $\theta$ is the angle between the direction of $\vec{L}_{s}$ and the $z$-axis, and $\varphi$ is the angle between the projection of $\vec{L}_{s}$ in the X-Y plane and the $x$-axis. There is no reason to assume the probability density $p$ depends on $\varphi$ since the external field is along the $z$-axis. We can simply drop the parameter $\varphi$ from $p(\theta,\varphi)$. The Larmor angular frequency is $\omega=eB_{z}/m$, which is independent of angle $\theta$. The corresponding potential energy for Larmor precession is
@@
Note that $d\varphi=\omega dt$ where $\varphi$ is the angle of circular rotation due to Larmor precession. In the infinitesimal period of time $\Delta t$, the electron intrinsic angular momentum rotates $\Delta\varphi=\omega\Delta t$. We can rewrite
@@
To compute $I_{f}$, we recall that the framework based on the extended least action principle Yang2023 allows us to choose general relative entropy definitions such as Kullback-Leibler divergence, Rényi divergence, or Tsallis divergence Tsallis ; Erven2014 ; Nielsen2011 . Here, we will choose Tsallis divergence for the reason to be clear later. For the infinitesimal period of time $\Delta t$, we define
@@
$\begin{split}A_{t}&=-\frac{1}{2}g_{s}L_{s}\int\int p(\theta)\cos(\theta)d\theta d\varphi\\ &+\frac{\hbar}{2(\alpha-1)}\{\int\frac{p^{\alpha}(\theta)}{\sigma^{\alpha-1}}d\theta d\varphi-1\}.\end{split}$ (9)
@@
where $Z_{\alpha}$ is a normalization factor that is dependent on $\alpha$. Now for $p(\theta)$ to be a probability density number, it must be real and non-negative. This imposes restrictions on the value of $\alpha$. Since $\cos(\theta)$ can be a negative number, $1/(\alpha-1)$ needs to be an even integer. Let $1/(\alpha-1)=2m$ where $m\in\mathbb{N}$. This gives
@@
Thus, from the extended least action principle and Assumption 3, we obtain a family of probability density functions $p_{m}(\theta)$. Mathematically, each of the probability density functions is a legitimate solution. We will further impose physical constraints in the next subsection and single out the solution that matches measurement results.
III.2 Discreteness of Spin Measurement
The probability density functions $p_{m}(\theta)$ in (13) are valid only relative to the context of magnetic field $B_{z}$ since $B_{z}$ defines the direction of the space to measure the spin. Without the external magnetic field, i.e., $B_{z}=0$, then the spin orientation is completely random. This corresponds to the case $m=0$ (or, $\alpha\to\infty)$ such that $p_{m}(\theta)$ is an uniform distribution. Recognizing the fact that spin measurement result is context dependent is itself an important result [^2] More accurately, we should label the probability density functions as $p_{m}[(\theta)|B_{z}]$ to show its contextual dependency. For simplicity of notation, we will not adopt such labeling in this paper. However, conceptual implication on such contextual dependency will be given in the Section VII..
When an inhomogeneous field $B_{z}$ is applied, $m$ must be a number larger than zero. This means $m$ increases when the spatial gradient of $B_{z}$ is non-zero. We can assume further that $m$ monotonically increases as the electron travels along the inhomogeneous field $B_{z}$. The justification of this assumption requires a physical model of electron. We will provide an intuitive explanation of this assumption and discuss possible physical model in Section VII.2. Here we just call out this assumption explicitly.
@@ III.2 Discreteness of Spin Measurement
Now we ask, what happens if the gradient of $B_{z}$ is sufficiently large, or the electron has been traveling in the inhomogeneous magnetic field long enough, such that $m$ can be approximated as infinity? Since $\theta\in[0,\pi]$ and $\cos(\theta)\in[-1,1]$, we have
@@
Thus, when $\nabla B_{z}$ is sufficiently large, the probability density function is
@@
This shows that when $\nabla B_{z}$ is sufficiently large, or the electron has been traveling in the inhomogeneous magnetic field for a distance long enough, measurement on the electron spin can only obtain two discrete outcomes: spin up and spin down, along the $z$-axis, as shown in Figure 2. Impressively, the quantum phenomenon of quantization in spin measurement outcome is a natural mathematical result here instead of being a postulate as usually taught in orthodox quantum mechanics textbooks.
However, (15) itself cannot be a valid probability density function. Its integral is zero since it is only non-zero when $\theta=0,\pi$. The Dirac delta function is the proper mathematical tool to describe such a scenario. We can rewrite
@@
The factor of $1/2$ is due to two facts, 1.) normalization requirement, and 2.) the measurement outcome of spin-up or spin-down is completely random. The second point here needs more elaboration. Before the external inhomogeneous field $B_{z}$ is applied, the intrinsic angular momentum of the electron is oriented completely randomly. At the moment when $B_{z}$ is applied, with half of chance that the initial direction of the angular momentum forms an angle with the $z$-axis $\theta$ such that $\theta\in[0,\pi/2]$. As the electron travels along $B_{z}$, the angular momentum vector evolves closer and closer to the $z$-axis and eventually becomes spin-up. With the other half of chance the initial direction of the angular momentum forms an angle with the $z$-axis $\theta$ such that $\theta\in[\pi/2,\pi]$. As the electron travels along $B_{z}$, the angular momentum evolves to be spin-down. For simpler notation, define
$\Theta^{+}:=\forall\theta\in[0,\pi/2],\mbox{ and }\Theta^{-}:=\forall\theta\in[\pi/2,\pi].$ (17)
Then, $\theta\in\Theta^{+}$ means the orientation is pointing to the upper half of the orientation sphere. Similarly, $\theta\in\Theta^{-}$ means the orientation is pointing to the lower half of the orientation sphere. The factor $1/2$ is thus due to initial complete randomness of orientation of intrinsic angular momentum such that the probability of $\theta\in\Theta^{+}$ is the same as the probability of $\theta\in\Theta^{-}$.
Figure 2: a.) As parameter $m$ increases, the probability distribution converges to only two directions at $\theta=0$ and $\theta=\pi$. b.) The envisioned Stern-Garlach measurements with an inhomogeneous magnetic field. As the electrons travel along $B_{z}$, the electrons reach the detectors that eventually show two converging discrete lines.
@@
We will derive (20) more rigorously in later sections. For an initially uniformly distributed orientation, one gets $p(\uparrow)=p(\downarrow)=1/2$.
The probability of $1/2$ for either spin-up and spin-down is due to the complete randomness of the initial orientations of the intrinsic angular momentum. It is important to note that the measurement results are relative to the direction defined by the magnetic field $B_{z}$. Given the same initial random orientations, if one measures the spin along a different direction, say $z^{\prime}$-axis, one will still get either spin-up and spin-down with probability of $1/2$. However, if the initial orientation of the intrinsic angular momentum is not completely random, the probability of obtaining spin-up and spin-down will be different from 1/2. This is the subject for study in the next section.
@@ III.3 Subsequent Measurement with a Rotated Stern-Gerlach Apparatus
Suppose after the electron passes through a Stern-Gerlach apparatus and it is observed the measurement result is spin-up. Now let the electron pass a subsequent Stern-Gerlach apparatus with the exact same direction of inhomogeneous magnetic field, one will obtain the result of spin-up for the electron. This means the electron stays in the spin-up state after passing through the first Stern-Gerlach apparatus [^3] In standard quantum mechanics, this is explained as that the electron stays in the eigen-state of spin-up.. However, in our model, after the electron passes through the first Stern-Gerlach apparatus and the external magnetic field is removed, due to the nature of random fluctuations, the orientation of electron intrinsic angular momentum can start to deviate from the $z$-axis with an angle $\theta$ between the angular momentum and the $z$-axis. In other words, due to fluctuations, the orientation of the angular momentum is relaxing away from pointing along the $z$-axis. This process is important in order to explain the behavior if the second Stern-Gerlach apparatus is set up such that the direction of the magnetic field is tilted with an angle $\beta$ from the original $z$-axis. Denote this new direction as $z^{\prime}$-axis, and the magnetic field as $B_{z^{\prime}}$.
Let the probability density conditioned on the initial measurement outcome of spin-up as $p(\theta^{\prime}|\theta)$, where $\theta^{\prime}$ is the angle between the orientation of electron intrinsic angular momentum and the $z^{\prime}$-axis. From Figure 3, it can be seen that $\theta=\theta^{\prime}+\beta$. We will apply the extended least action principle again to find out $p(\theta^{\prime}|\theta)$. With the magnetic field $B_{z^{\prime}}$, the classical action can be calculated similarly to that in (6) - (7) as
@@
To compute $I_{f}$, we choose Tsallis divergence again. For the infinitesimal period of time $\Delta t$, we define
@@
Note that (29) is the same as (20) if we set $\varrho(\beta)=\varrho(\Theta^{+})$ and $\varrho(\beta+\pi)=\varrho(\Theta^{-})$. This effectively proves (20).
The average of angular momentum along the $z^{\prime}$-axis will be $(\varrho(\beta)-\varrho(\beta+\pi))L_{s}$. On the other hand, from Figure 3, one can deduce that this average angular momentum must be $L_{s}\cos(\beta)$, thus [^4] If the initial condition is spin-down after the electron passing the first Stern-Gerlach apparatus, the average angular momentum becomes $L_{s}\cos(\beta)$, resulting $\varrho(\beta)-\varrho(\beta+\pi)=-\cos(\beta)$.
@@
Thus, we reproduce the same result as that from the standard quantum mechanics,
@@
Suppose $\beta<\pi/2$, we wonder why there is possibility that the measurement result of the second Stern-Gerlach apparatus is spin-down. That is, why $p(\downarrow|\beta)$ is non-zero? The question arises because if the initial condition is spin-up after the electron passes through the first Stern-Gerlach apparatus, and if the orientation of the intrinsic angular momentum is kept unchanged, the initial angle $\theta^{\prime}=\beta\in(0,\pi/2)$. Then, according to the discussions leading to (18), we should only obtain spin-up as the only results when $B_{z^{\prime}}$ is applied. However, as pointed out at the beginning of this section, after the electron passing through the first Stern-Gerlach apparatus, the direction of the intrinsic angular momentum starts to deviate from the $z$-axis due to vacuum fluctuations. Suppose after a sufficient time of relaxation $\tau_{r}$, the orientation of electron intrinsic angular momentum can be at any angle $\theta$ from the $z$-axis with $\theta\in[0,\pi/2]$. When the electron enters the second Stern-Gerlach with the magnetic field $B_{z^{\prime}}$ is tilted with an angle $\beta\in(0,\pi/2)$, the initial orientation is now measured with respect to the $z^{\prime}$-axis as angle $\theta^{\prime}$. Since $\theta^{\prime}=\theta-\beta$, the orientation of the intrinsic angular momentum can point to either upward with $\theta^{\prime}\in(\Theta^{\prime})^{+}$, or downward with $\theta^{\prime}\in(\Theta^{\prime})^{-}$, as shown in Figure 3b. Thus, there is possibility of spin-down when $\nabla B_{z^{\prime}}$ is sufficiently large. We see the relaxation of intrinsic angular momentum orientation is necessary in order to explain the result in (32), which has been confirmed experimentally.
There is still one puzzle to clarify. The second Stern-Gerlach apparatus can be rotated such that $\theta-\theta^{\prime}$ is $\beta$ or $\beta$, the probability of measurement outcome for spin-up or spin-down (33) are the same for both cases. This suggests that (33) does not completely specify the behavior of electron spin in a single measurement. To remedy that, we need to specify the probability of measurement outcome along another axis, say, the $y$-axis. Suppose the yzy-z plane is rotated along the $x$-axis such that $z^{\prime}$-axis is deviated from $z$-axis by an angle $\beta$, then the angle between $y^{\prime}$-axis and $z$-axis is $\pi/2+\beta$. The probability density of measurement outcome along the $y^{\prime}$-axis can be obtained by replacing parameter $\beta$ in (33) with $\pi/2+\beta$,
@@
On the other hand, if the $z^{\prime}$-axis is deviated from $z$-axis by an angle $\beta$, then the angle between $y^{\prime}$-axis and $z$-axis is $\pi/2-\beta$. The probability density of measurement outcome along the $y^{\prime}$-axis is
@@
It is clear that (34) and (35) are different. Thus, (33) and (34) together give a complete description of spin behavior when then second Stern-Gerlach is rotated by an angle $\beta$, while (33) and (35) together give a complete description of spin behavior when then second Stern-Gerlach is rotated by an angle $\beta$.
We can generalize the result by assuming the magnetic field of the first Stern-Gerlach apparatus is along a direction tilted an angle $\beta_{1}$ from the $z$-axis, and the second Stern-Gerlach apparatus is along a direction tilted an angle $\beta_{2}$ from the $z$-axis. Supposed measurement outcome of the first Stern-Gerlach apparatus is spin-up, what is the probability of measurement outcome with spin-up from the second Stern-Gerlach apparatus? The calculation steps from (21) to (32) can be repeated but with $\theta=\theta^{\prime}+\beta_{2}-\beta_{1}$. Thus, we can just replace $\beta$ with $\beta_{2}-\beta_{1}$ in steps from (21) to (32), and the final result will be

### v3

@@ III.1 Spin Model
In order to explain the existence of electron spin, additional assumptions on electron properties are needed. Historically, several models have been proposed to explain the existence of spin. The most popular model is to consider the electron as a rotating rigid body with uniform distributed charge. Denote the magnetic moment of the electron as $\vec{\mu}$ and the angular momentum as $\vec{L}$. The relation between $\vec{\mu}$ and $\vec{L}$ can be derived if we consider the electron as a localized electrical distribution with current density $\vec{j}(\vec{r})$, and the magnetic moment of electron is the magnetic dipole moment, which is calculated as
@@
where $\vec{p}(\vec{r})$ is the momentum density. However, one of the problems of this model is that when the radius of the rigid body is small enough, the edge moves faster than the speed of light Tomonaga . There is also a simpler model that considers an electron as a point particle with intrinsic magnetic moment and intrinsic angular momentum, and they still satisfy the relation (3). Such a model does not offer an explanation as to how the intrinsic angular moment or the intrinsic magnetic moment originates.
Based on the success of applying the extended least action principle in deriving quantum theories Yang2023 ; Yang2024 , we wish to explain the origin of the intrinsic magnetic moment and angular momentum by applying the same principle but adding additional refinement on the assumption of the vacuum fluctuations in Section II. The electron exhibits random displacement as a result of vacuum fluctuation, and this displacement is described as a vector. The displacement vector is intrinsically local in the sense that it is independent of the reference of origin of the coordinate system. If we further assume that the displacement vector contains not only translational components, but also rotational components, then the electron will exhibit intrinsic angular momentum as a result of such random rotation. In Appendix B, using the extended least action principle, we show that if the rotation component of the displacement vector is a circular movement, the averaged magnitude of intrinsic angular momentum turns out to be $L_{s}=\hbar/2$, while the orientations of intrinsic angular momentum are completely random. That is, the intrinsic angular momentum is isotropic and its orientation randomly fluctuates. Although our model is very different from the rigid body model, we assume that the relation (3) still holds except with a factor $g_{s}$ to be determined further
@@
Since the orientation of the magnetic moment is always opposite to the orientation of the intrinsic angular momentum, the orientation of the intrinsic magnetic moment fluctuates randomly as well. Although the result in Appendix B is impressive enough, the assumption of circular motion of displacement vectors due to vacuum fluctuation is rather strong [^1] It is desirable to develop a better model than that is described in Appendix B to derive the $\hbar/2$ magnitude of intrinsic angular momentum, but we leave it as a future research topic.. Instead, we will just assume the existence of an intrinsic angular momentum of magnitude $\hbar/2$ and with a random orientation for an electron. In summary, in addition to Assumptions 1 and 2 as described in Section II, we need
@@
This additional assumption allows us to apply the extended least action principle to derive the probability density of the orientation of the intrinsic angular momentum.
Suppose that we choose a reference frame such that the electron is at rest, that is, the average translational momentum is zero. The electron still exhibits an intrinsic angular momentum with random orientation according to Assumption 3. We want to derive the probability distribution of the intrinsic angular momentum orientations. If no external magnetic field is applied, the probability distribution is simply a uniform distribution according to Assumption 3. Now, suppose that an external magnetic field along the direction of the $z$-axis, $B_{z}$, is applied, the probability distribution of the magnetic moment orientations is no longer uniform. We show next that this probability distribution can be derived from the extended least action principle.
Due to the interaction of the electron magnetic moment and the external magnetic field, the electron is experiencing Larmor precession around the $z$-axis, as shown in Figure 1. Denote the probability density of the intrinsic angular momentum orientations as $p(\theta,\varphi)$ where $\theta$ is the angle between the direction of $\vec{L}_{s}$ and the $z$-axis, and $\varphi$ is the angle between the projection of $\vec{L}_{s}$ in the X-Y plane and the $x$-axis. There is no reason to assume that the probability density $p$ depends on $\varphi$ since the external field is along the $z$-axis. We can simply drop the parameter $\varphi$ from $p(\theta,\varphi)$. The Larmor angular frequency is $\omega=eB_{z}/m$, which is independent of angle $\theta$. The corresponding potential energy for Larmor precession is
@@
Note that $d\varphi=\omega dt$ where $\varphi$ is the angle of circular rotation due to Larmor precession. In an infinitesimal period of time $\Delta t$, the intrinsic angular momentum of the electron rotates $\Delta\varphi=\omega\Delta t$. We can rewrite
@@
To compute $I_{f}$, we recall that the framework based on the extended least action principle Yang2023 allows us to choose a general relative entropy definition such as Kullback-Leibler divergence, Rényi divergence, or Tsallis divergence Tsallis ; Erven2014 ; Nielsen2011 . The Rényi divergence, or the Tsallis divergence, is the one-parameter generation of the K-L divergence, where the parameter is called the order of divergence. The K-L divergence is a special limit when the order of divergence approaches one. As will be seen later, the parameter of the order of divergence is needed to explain spin quantization. Therefore, we exclude the K-L divergence and can choose either the Rényi divergence or the Tsallis divergence. There is no fundamental reason to choose one or the other, although the Tsallis divergence has the advantage of mathematical simplicity. In Appendix C, we prove that choosing Rényi divergence gives the similar results on spin quantization. With these considerations, we define $I_{f}$ for the infinitesimal period of time $\Delta t$ as follows.
@@
$\begin{split}A_{t}&=-\frac{1}{2}g_{s}L_{s}\int p(\theta)\cos(\theta)d\theta d\varphi\\ &+\frac{\hbar}{2(\alpha-1)}\{\int\frac{p^{\alpha}(\theta)}{\sigma^{\alpha-1}}d\theta d\varphi-1\}.\end{split}$ (9)
@@
where $Z_{\alpha}$ is a normalization factor that is dependent on $\alpha$. Now, for $p(\theta)$ to be a probability density number, it must be real and non-negative. This imposes restrictions on the value of $\alpha$. Since $\cos(\theta)$ can be a negative number, $1/(\alpha-1)$ should be an even integer. Let $1/(\alpha-1)=2m$ where $m\in\mathbb{N}$. This gives
@@
Thus, from the extended least action principle and Assumption 3, we obtain a family of probability density functions $p_{m}(\theta)$. Mathematically, each of the probability density functions is a legitimate solution. We will further impose physical constraints in the next subsection and single out the solution that can explain the measurement results of the electron spin.
III.2 Discreteness of Spin Measurement Results
The probability density functions $p_{m}(\theta)$ in (13) are valid only relative to the context of the magnetic field $B_{z}$ since $B_{z}$ defines the direction of the space to measure the spin. Without the external magnetic field, i.e., $B_{z}=0$, then the spin orientation is completely random. This corresponds to the case $m=0$ (or $\alpha\to\infty)$ such that $p_{m}(\theta)$ is a uniform distribution. Recognizing the fact that spin measurement result is context dependent is itself an important result [^2] More precisely, we should label the probability density functions as $p_{m}[(\theta)|B_{z}]$ to show its contextual dependency. For simplicity of notation, we will not adopt such labeling in this paper. However, the conceptual implications on this contextual dependency will be discussed in Section VII..
When an inhomogeneous field $B_{z}$ is applied in the Stern-Gerlach experiment, the electron movement is no longer randomly distributed along the $z$ -axis. This means that we need to choose $p_{m}(\theta)$ with $m$ larger than zero. However, there are infinite numbers of $m>0$ to choose from. We only know that the value of $m$ depends on the inhomogeneous field $B_{z}$. The exact relation between the value of $m$ and the inhomogeneous field $B_{z}$ cannot be derived from the extended least action principle. At this point, we need to make another assumption on the relation between the value of $m$ and the inhomogeneous field $B_{z}$, that $m$ monotonically increases as the electron travels along the inhomogeneous field $B_{z}$. The justification of this assumption requires a more detailed physical model of electron. However, we will provide an intuitive explanation of this assumption and a possible physical model in Section VII.2. The intention here is to make minimal assumptions so that we can recover the quantum theory of the electron spin, including the discreteness of spin measurement results, the Pauli-Schrodinger equation, and entanglement between two correlated electron spins, as shown in later sections. Explicitly, we have
@@ III.2 Discreteness of Spin Measurement Results
With this assumption, one can extrapolate the outcomes if the gradient of $B_{z}$ is sufficiently large, or the electron has been traveling in the inhomogeneous magnetic field long enough such that $m$ can be approximated as infinity. Since $\theta\in[0,\pi]$ and $\cos(\theta)\in[-1,1]$, we have
@@
Thus, the probability density function is
@@
This shows that when $\nabla B_{z}$ is sufficiently large, or the electron has been traveling in the inhomogeneous magnetic field for a distance long enough, measurement of the electron spin can only obtain two discrete outcomes: spin up and spin down, along the $z$-axis, as shown in Figure 2. This explains the quantization of spin measurement outcomes.
However, (15) itself cannot be a valid probability density function. Its integral is zero since it is only nonzero when $\theta=0,\pi$. The Dirac delta function is the proper mathematical tool to describe such a situation. We can rewrite
@@
The factor of $1/2$ is due to two facts, 1.) the normalization requirement, and 2.) the result of the spin-up or spin-down measurement is completely random. The second point here needs more elaboration. Before applying the external inhomogeneous field $B_{z}$, the intrinsic angular momentum of the electron is completely randomly oriented. At the moment when $B_{z}$ is applied, it is half the chance that the initial direction of the angular momentum forms an angle with the $z$-axis $\theta$ such that $\theta\in[0,\pi/2]$. As the electron travels along $B_{z}$, the angular momentum vector evolves closer and closer to the $z$ -axis and eventually spins up. With the other half of chance, the initial direction of the angular momentum forms an angle with the $z$-axis $\theta$ such that $\theta\in[\pi/2,\pi]$. As the electron travels along $B_{z}$, the angular momentum evolves to be spin-down. For simpler notation, define
$\Theta^{+}:=\{\forall\theta\in[0,\pi/2]\},\mbox{ and }\Theta^{-}:=\{\forall\theta\in[\pi/2,\pi]\}.$ (17)
Then, $\theta\in\Theta^{+}$ means that the spin orientation is pointing to the upper half of the orientation sphere. Similarly, $\theta\in\Theta^{-}$ means that the spin orientation is pointing to the lower half of the orientation sphere. The factor $1/2$ is therefore due to the complete initial randomness of the orientation of the intrinsic angular momentum such that the probability of $\theta\in\Theta^{+}$ is the same as the probability of $\theta\in\Theta^{-}$.
Figure 2: a.) As parameter $m$ increases, the probability distribution converges to only two directions at $\theta=0$ and $\theta=\pi$. b.) The envisioned Stern-Garlach measurement results with an inhomogeneous magnetic field. As the electrons travel along $B_{z}$, the electrons reach the detectors that eventually show two converging discrete lines.
@@
We will derive (20) more rigorously in later sections. For an initially uniformly distributed orientation, $\sigma$ is a constant, and we get $p(\uparrow)=p(\downarrow)=1/2$.
The probability of $1/2$ for spin-up and spin-down is due to the complete randomness of the initial orientations of the intrinsic angular momentum. It is important to note that the measurement results are relative to the direction defined by the magnetic field $B_{z}$. Given the same initial random orientations, if one measures the spin along a different direction, say $z^{\prime}$-axis, one will still get either spin-up and spin-down with probability of $1/2$. However, if the initial orientation of the intrinsic angular momentum is not completely random, the probability of obtaining spin-up and spin-down along the $z^{\prime}$-axis will be different from 1/2. This is the subject of study in the next section.
@@ III.3 Subsequent Measurement with a Rotated Stern-Gerlach Apparatus
Suppose after the electron passes through a Stern-Gerlach apparatus and is observed the measurement result is spin-up. Now, letting the electron pass a subsequent Stern-Gerlach apparatus with the exact same direction of inhomogeneous magnetic field, one will obtain the result of spin-up for the electron. This means that the electron remains in the spin-up state after passing through the first Stern-Gerlach apparatus [^3] In standard quantum mechanics, this is explained as that the electron stays in the eigen-state of spin-up.. However, in our model, after the electron passes through the first Stern-Gerlach apparatus and the external magnetic field is removed, due to the nature of random fluctuations, the orientation of the intrinsic electron angular momentum can start to deviate from the $z$-axis with an angle $\theta$ between the angular momentum and the $z$-axis. In other words, due to fluctuations, the orientation of the angular momentum is relaxing away from pointing along the $z$-axis. This process is important in order to explain the behavior if the second Stern-Gerlach apparatus is set up such that the direction of the magnetic field is tilted with an angle $\beta$ from the original $z$-axis. Denote this new direction by $z^{\prime}$-axis and the magnetic field by $B_{z^{\prime}}$.
Let the probability density conditioned on the initial measurement outcome of spin-up be $p(\theta^{\prime}|\theta)$, where $\theta^{\prime}$ is the angle between the orientation of the intrinsic angular momentum of the electron and the $z^{\prime}$-axis. From Figure 3, it can be seen that $\theta=\theta^{\prime}+\beta$. We will apply the extended least action principle again to find out $p(\theta^{\prime}|\theta)$. With the magnetic field $B_{z^{\prime}}$, the classical action can be calculated similarly to that in (6) - (7) as
@@
To compute $I_{f}$, we choose the Tsallis divergence again. For the infinitesimal period of time $\Delta t$, we define
@@
Notice that (29) is the same as (20) if we set $\varrho(\beta)=\varrho(\Theta^{+})$ and $\varrho(\beta+\pi)=\varrho(\Theta^{-})$. This effectively proves (20).
The average angular momentum along the $z^{\prime}$-axis will be $(\varrho(\beta)-\varrho(\beta+\pi))L_{s}$. On the other hand, from Figure 3, one can deduce that this average angular momentum must be $L_{s}\cos(\beta)$, thus [^4] If the initial condition is spin-down after the electron passing the first Stern-Gerlach apparatus, the average angular momentum becomes $L_{s}\cos(\beta)$, resulting $\varrho(\beta)-\varrho(\beta+\pi)=-\cos(\beta)$.
@@
Therefore, we reproduce the same result as that from the standard quantum mechanics,
@@
Suppose that $\beta<\pi/2$. We wonder why there is the possibility that the measurement result of the second Stern-Gerlach apparatus is spin-down. That is, why $p(\downarrow|\beta)$ is non-zero? The question arises because if the initial condition is spin-up after the electron passes through the first Stern-Gerlach apparatus, and if the orientation of the intrinsic angular momentum is kept unchanged, the initial angle $\theta^{\prime}=\beta\in(0,\pi/2)$. Then, according to the discussions leading to (18), we should only obtain spin-up as the only result when $B_{z^{\prime}}$ is applied. However, as pointed out at the beginning of this section, after the electron passes through the first Stern-Gerlach apparatus, the direction of the intrinsic angular momentum starts to deviate from the $z$-axis due to vacuum fluctuations. Suppose after a sufficient relaxation time $\tau_{r}$, the orientation of the intrinsic angular momentum of the electron can be at any angle $\theta$ from the $z$-axis with $\theta\in[0,\pi/2]$. When the electron enters the second Stern-Gerlach with the magnetic field $B_{z^{\prime}}$ being tilted at an angle $\beta\in(0,\pi/2)$, the initial orientation is now measured with respect to the $z^{\prime}$-axis as angle $\theta^{\prime}$. Since $\theta^{\prime}=\theta-\beta$, the orientation of the intrinsic angular momentum can point to either upward with $\theta^{\prime}\in(\Theta^{\prime})^{+}$, or downward with $\theta^{\prime}\in(\Theta^{\prime})^{-}$, as shown in Figure 3b. Thus, there is a possibility of spin-down when $\nabla B_{z^{\prime}}$ is sufficiently large. We see that relaxation of the orientation of intrinsic angular momentum is necessary to explain the result in (32), which has been confirmed experimentally.
There is still a puzzle to be clarified. The second Stern-Gerlach apparatus can be rotated so that $\theta-\theta^{\prime}$ is $\beta$ or $\beta$, the probability of measurement outcome for spin-up or spin-down (33) is the same for both cases. This suggests that (33) does not completely specify the behavior of the electron spin in a single measurement. To remedy that, we need to specify the probability of measurement outcome along another axis, say the $y$-axis. Suppose the yzy-z plane is rotated along the $x$-axis such that $z^{\prime}$-axis is deviated from $z$-axis by an angle $\beta$, then the angle between $y^{\prime}$-axis and $z$-axis is $\pi/2+\beta$. The probability density of measurement outcome along the $y^{\prime}$-axis can be obtained by replacing parameter $\beta$ in (33) with $\pi/2+\beta$,
@@
On the other hand, if the $z^{\prime}$-axis is deviated from the $z$-axis by an angle $\beta$, then the angle between the $y^{\prime}$-axis and the $z$-axis is $\pi/2-\beta$. The probability density of measurement outcome along the $y^{\prime}$-axis is
@@
It is clear that (34) and (35) are different. Thus, (33) and (34) together give a complete description of spin behavior when the second Stern-Gerlach is rotated by an angle $\beta$, while (33) and (35) together give a complete description of spin behavior when the second Stern-Gerlach is rotated by an angle $\beta$.
We can generalize (32) by assuming that the magnetic field of the first Stern-Gerlach apparatus is along a direction tilted at an angle $\beta_{1}$ from the $z$-axis, and the second Stern-Gerlach apparatus is along a direction tilted at an angle $\beta_{2}$ from the $z$-axis. Suppose that measurement outcome of the first Stern-Gerlach apparatus is spin-up, what is the probability of measurement outcome with spin-up from the second Stern-Gerlach apparatus? The calculation steps from (21) to (32) can be repeated but with $\theta=\theta^{\prime}+\beta_{2}-\beta_{1}$. Thus, we can just replace $\beta$ with $\beta_{2}-\beta_{1}$ in steps from (21) to (32), and the final result will be